A stock had returns of 18.55 percent, −7.75 percent, and 24.05 percent for the past three years. What is the standard deviation of the returns?
Answer #1

First we compute the average returnsas (18.55%-7.75%+24.05%)/3 = 11.6167%

Next calculate sum of (Return- Averge )^2

= (18.55%-11.6167%)^2+(-7.75%-11.6167%)^2+ (24.05-11.6167%)^2

= 577.7267

Standard deviation = SQRT( sum of (Return- Averge )^2/ (n-1))

= SQRT(577.7267/2)

=

16.99598 %

=17%